Unpublished talks without teaching courses Hans–Paul Schwefel

August 3, 2022

[1] Hans-Paul Schwefel. Wandschubspannungsmessungen bei ablösender und abgelöster Grenzschicht. Teestunden-Seminar des Hermann-Föttinger-Instituts für Strömungstechnik, Technische Universität (TU) Berlin, June 19, 1964.

[2] Hans-Paul Schwefel. Kybernetische Lösungsansteuerung experimenteller Optimierungsprobleme. Gemeinsames Seminar der Projektgruppe MHD-Staustrahlrohr des AEG-Forschungsinstituts Berlin und des Instituts für Kerntechnik der TU Berlin, April 1, 1966.

[3] Hans-Paul Schwefel. Das Studium an der TU Berlin. Abiturienten-Beratung am Canisius-Kolleg Gymnasium, Berlin-Tiergarten, October 24, 1966.

[4] Hans-Paul Schwefel. Vergleich zwischen Evolutions- und Gradientenstrategie. Fortgeschrittenenstudium Adaptive und lernende Systeme, Brennpunkt Kybernetik an der TU Berlin, November 18, 1966.

[5] Hans-Paul Schwefel. Studium und Berufsaussichten des Diplomingenieurs. Abiturienten-Beratung am Canisius-Kolleg Gymnasium, Berlin-Tiergarten, September 14, 1968.

[6] Peter Bienert, Ingo Rechenberg, and Hans-Paul Schwefel. Evolutionsstrategie. Industrieausstellung Berlin, Sonderschau Wissenschaftler in Gruben, Stand Biotechnik zusammen mit Johann–Gerhard Helmcke und Frei Otto, September 27 – October 6, 1969.

[7] Hans-Paul Schwefel. Wärmeübertragungsanlage für Flüssigmetall bis 900 Grad Celsius. Seminar des AEG-Forschungsinstituts, Frankfurt/Main, October 16, 1969.

[8] Hans-Paul Schwefel. Stand der Arbeiten beim Projekt MHD-Staustrahlrohr. Seminar der Arbeitsgruppe II Antriebe und Energiequellen der Kommission für Raumfahrt (KfR), Braunschweig, November 13, 1969.

[9] Hans-Paul Schwefel. Künstliche Evolution – eine Strategie der experimentellen Optimierung. Seminar des AEG-Forschungsinstituts, Berlin, December 2, 1969.

[10] Hans-Paul Schwefel. Künstliche Evolution – eine Strategie der experimentellen Optimierung. Kolloquium des Instituts für Raumfahrt (Prof. Ruppe), Technische Hochschule Munich, December 4 1969.

[11] Hans-Paul Schwefel. Künstliche Evolution – eine Strategie der experimentellen Optimierung. VDE Niederbayern, Amberg/Oberpfalz, January 14, 1970.

[12] Hans-Paul Schwefel. Studium und Berufsaussichten des Diplomingenieurs. Abiturienten-Beratung am Canisius-Kolleg Gymnasium, Berlin-Tiergarten, July 1, 1970.

[13] Hans-Paul Schwefel. Künstliche Evolution – eine Strategie der experimentellen Optimierung. VDI-Arbeitskreis Bauingenieurwesen, Dusseldorf, December 17, 1970.

[14] Peter Bienert, Ingo Rechenberg, and Hans-Paul Schwefel. Evolutionstechnik. Informationsseminar der Firma Messerschmidt-Bölkow-Blohm (MBB), Ottobrunn, July 6–8, 1971.

[15] Hans-Paul Schwefel. Evolutionsstrategien. Deutsches Institut für Entwicklungspolitik, Berlin, October 11, 1971.

[16] Hans-Paul Schwefel. Optimierverfahren und ihre Anwendungsmöglichkeiten. Deutsches Institut für Entwicklungspolitik, Berlin, November 1, 1971.

[17] Hans-Paul Schwefel. Zufall und Notwendigkeit – über ein Buch von Jacques Monod. Pfarrer-Wiesinger-Kreis, Berlin-Borsigwalde, November 25, 1971.

[18] Hans-Paul Schwefel. Methoden der numerischen Optimierung. Seminar des Instituts für Mess- und Regelungstechnik, TU Berlin, January 31, 1972.

[19] Peter Bienert, Ingo Rechenberg, and Hans-Paul Schwefel. Optimierung mit der Evolutionsstrategie. Fortbildungsseminar Automatisierung – Analyse und Synthese dynamischer Systeme, Brennpunkt Kybernetik an der TU Berlin, October 8–9, 1973.

[20] Hans-Paul Schwefel. Ein Vergleich direkter Methoden der Parameteroptimierung. Arbeitskreis Optimierprogramme der Dechema, Frankfurt/Main, October 24, 1974.

[21] Hans-Paul Schwefel. Erfahrungen mit der Evolutionsstrategie. Arbeitskreis Optimierprogramme der Dechema, Frankfurt/Main, April 10, 1975.

[22] Hans-Paul Schwefel. Zuverlässigkeit und Aufwand der Evolutionsstrategie – ein Vergleich mit den wichtigsten direkten Suchmethoden der Optimierung. Seminar am Wissenschaftlichen Zentrum der IBM, Heidelberg, April 18, 1975.

[23] Hans-Paul Schwefel. Direkte Suchstrategien zur Parameteroptimierung. Seminar der Programmgruppe Systemforschung und Technologische Entwicklung (STE), Kernforschungsanlage Jülich (KFA), May 25, 1976.

[24] Hans-Paul Schwefel. Entwicklung der Energiewirtschaft in der Bundesrepublik – Ergebnisse von Modellrechnungen. Seminar des Instituts für Reaktorbauelemente, Kernforschungszentrum Karlsruhe (KFK), April 20, 1977.

[25] Hans-Paul Schwefel. Optimierung von Simulationsmodellen mit der Evolutionsstrategie. Workshop Simulationsmethoden in der Medizin und Biologie, Medizinische Hochschule Hanover, September 29, 1977.

[26] Hans-Paul Schwefel. Energy Modelling Activities at KFA-STE. Seminary at the Energy Study Centre, Netherlands Energy Research Foundation (ECN), Petten, The Netherlands, October 5, 1977.

[27] Hans-Paul Schwefel. Systemanalyse der Energieversorgung. Ferienkurs Röntgen- und Neutronenstreuung in der Festkörperphysik, Institut für Festkörperforschung (IFF) der KFA Jülich, April 7, 1978.

[28] Hans-Paul Schwefel. Evolutionsstrategie mit variabler Metrik. Fachseminar Ausgewählte Kapitel der Bionik und Evolutionstechnik, Institut für Mess- und Regelungstechnik, Fachbereich Verfahrenstechnik, TU Berlin, June 19, 1978.

[29] Hans-Paul Schwefel. Direct Search for Optimal Parameters Within Simulation Models (poster presentation). Symposium on Modelling and Simulation Methodology, The Weizmann Institute of Science, Rehovot, Israel, August 13–18, 1978.

[30] Hans-Paul Schwefel. Probleme bei der Anwendung der Evolutionsstrategie. Fachseminar Ausgewählte Kapitel der Bionik und Evolutionstechnik, Institut für Mess- und Regelungstechnik, Fachbereich Verfahrenstechnik, TU Berlin, February 1, 1979.

[31] Hans-Paul Schwefel. Energy Modelling Activities at KFA-STE. Boris Kidric Institute of Nuclear Sciences, Vinca, Belgrade, Yugoslavia, October 31, 1979.

[32] Hans-Paul Schwefel. Energy Systems Management by Computer Aided Analysis and Synthesis. IIASA-IFAC Symposium on Modeling of Large-scale Energy Systems, International Institute for Applied Systems Analysis (IIASA), Laxenburg, Austria, February 25–29, 1980.

[33] Hans-Paul Schwefel. Optimierung nach dem Vorbild der Evolution. VDI-Arbeitskreis der Konstruktionsingenieure (ADKI), Fachhochschule Bielefeld, April 24, 1980.

[34] Hans-Paul Schwefel. A National Energy Model for the F.R.G. and Its Software Environment. Vingtdeuxiéme Congrés National de la Société Canadienne pour la Recherche Opérationelle (SCRO), Québec, Canada, May 20–22, 1980.

[35] Hans-Paul Schwefel. Evolutionsstrategien. Seminar des Fachbereichs Luft- und Raumfahrttechnik, Hochschule der Bundeswehr, Neubiberg, June 2, 1980.

[36] Hans-Paul Schwefel. Evolutionsmechanismen aus der Sicht der numerischen Optimierung. Werkstattgespräch Evolution der Evolutionsmechanismen, Sonderforschungsbereich Versuchstierforschung (SFB 146), Hannover, und Max-Planck-Institut für biophysikalische Chemie, Göttingen, Schloss Berlepsch, October 31 – November 1, 1980.

[37] Hans-Paul Schwefel. Optimierung technischer Systeme mit der Evolutionsstrategie. Seminar der Österreichischen Gesellschaft für Operations Research (ÖGOR) und der Österreichischen Computer Gesellschaft (OCG), Vienna University of Technology, Austria, November 17, 1980.

[38] Friedhelm Drepper, Rainer Heckler, and Hans-Paul Schwefel. Langfristziele der Abteilung Rechnergestützte Planungsmethoden (REPLAN). Arbeitsseminar der Programmgruppe STE, KFA Jülich, December 2–5, 1980.

[39] Hans-Paul Schwefel. Die Organisationsstruktur der KFA. Seminar der Programmgruppe STE, KFA Jülich, January 20, 1981.

[40] Hans-Paul Schwefel. Global 2000 – der Bericht an den Präsidenten. Seminar der Programmgruppe STE, KFA Jülich, February 10, 1981.

[41] Hans-Paul Schwefel. Experimentelle und numerische Optimierung mit der Evolutionsstrategie. Seminar der Abteilung Systemtheorie, Institut für Systemwissenschaften, Johannes-Kepler-Universität, Linz, Austria, March 19, 1981.

[42] Hans-Paul Schwefel. Evolutionsstrategie – ein der Natur abgeschautes Optimierungskonzept. Außeninstitut der Technischen Universität Graz, Österreichischer Verband der Wirtschaftsingenieure, Graz, Austria, May 14, 1981.

[43] Hans-Paul Schwefel. Zur Nachahmung von Evolutionsprinzipien beim experimentellen und numerischen Optimieren. Kolloquium des Fachbereichs Informatik, Universität Dortmund, May 26, 1981.

[44] Hans-Paul Schwefel. Global 2000 – der Bericht an den Präsidenten. Aussagen, Methodik und erste Reaktionen. Fortbildungsseminar der KFA J”ulich, July 7, 1981.

[45] Friedhelm Drepper, Rainer Heckler, Ulrich Hermes, and Hans-Paul Schwefel. Software Exhibition: DAIMOS, IRECA, and OASIS. Ninth IFORS Conference Operational Research’81, Hamburg, July 21, 1981.

[46] Hans-Paul Schwefel. Evolutionsstrategien – ein neues Optimierungskonzept; mit einem Leistungsvergleich ableitungsfreier Verfahren der Parameteroptimierung. Seminar Mathematische Modelle und Methoden, Institut für Datenverarbeitung in der Technik, Kernforschungszentrum Karlsruhe (KFK), September 17, 1981.

[47] Hans-Paul Schwefel. Systemanalyse – Versuch einer Bestandsaufnahme. Kolloquium des Fachbereichs Informatik, Universität Dortmund, September 25, 1981.

[48] Hans-Paul Schwefel. Anwendung von Evolutionsstrategien zur Optimierung des Kraftwerkseinsatzes. Jahrestagung der Österreichischen Gesellschaft für Operations Research (ÖGOR), Vienna, Austria, September 28–30, 1981.

[49] Hans-Paul Schwefel. Das Energiebilanz-Auskunftssystem EBAS. Seminar der Programmgruppe STE, KFA Jülich, October 13, 1981.

[50] Hans-Paul Schwefel. Ein Zweisektoren-Modell, Stabilit”at und Evolutionsstrategien. Seminar der Programmgruppe STE, KFA Jülich, April 16, 1982.

[51] Friedhelm Drepper and Hans-Paul Schwefel. Übergang vom Gleichgewicht zum Chaos bei einem einfachen Produktionssystem mit Investitionsverflechtungen. Ferienkurs Nichtlineare Dynamik in kondensierter Materie, Institut für Festkörperforschung (IFF), KFA Jülich, March 11, 1983.

[52] Hans-Paul Schwefel. Suchstrategien in den Geistes- und Naturwissenschaften und die Möglichkeit zur prospektiven Abschätzung von Forschungsprogrammen oder: Determinismus versus Zufall als Forschungsstrategie. 23. Sitzung der Senatskommission für Versuchstierforschung der Deutschen Forschungsgemeinschaft (DFG), Mittelkalbach/Fulda, July 1–2, 1983.

[53] Hans-Paul Schwefel. Summary Report on Session 4 on Systems Analytical Approaches to International Stability (as chairman). IFAC Workshop Supplemental Ways for Improving International Stability (SWIIS), Vienna, Austria, September 13–15, 1983.

[54] Hans-Paul Schwefel. Ein Software-Konzept zur Analyse und Synthese komplexer Systeme. Institut für Systemwissenschaften, Technisch-Naturwissenschaftliche Fakultät, Johannes-Kepler-Universität, Linz, Austria, October 20, 1983.

[55] Friedhelm Drepper, Rainer Heckler, Ulrich Hermes, and Hans-Paul Schwefel. Software für die Analyse dynamischer Systeme – Teil 1. Workshop der Programmgruppe STE, KFA Jülich, March 14, 1984.

[56] Friedhelm Drepper, Rainer Heckler, Ulrich Hermes, and Hans-Paul Schwefel. Software für die Analyse dynamischer Systeme – Teil 2. Workshop der Programmgruppe STE, KFA Jülich, March 29, 1984.

[57] Hans-Paul Schwefel. Evolution und Optimierung – über ein paar experimentelle und numerische Versuche. Max Planck Institute for Biophysical Chemistry, Göttingen, October 2, 1985.

[58] Hans-Paul Schwefel. Evolutionsstrategien. Interdisciplinary Working Group of Mathematization, University of Kassel, February 6, 1986.

[59] Hans-Paul Schwefel. Evolution strategies: A family of nonlinear optimization techniques based on imitating some principles of organic evolution. Institute of Systems Research of the Polish Academy of Science, Warsaw, August 20, 1986.

[60] Hans-Paul Schwefel. Weltmodelle. Siemens AG Munich, Central Department of Research and Technology, March 25, 1987.

[61] Hans-Paul Schwefel. Evolutionsstrategien in der Optimierungstheorie. Projektgruppe Simulation evolutionärer Prozesse, University of Dortmund, Department of Mathematics, April 24, 1987.

[62] Hans-Paul Schwefel. Evolution und Optimierung – Evolutionsstrategie(n). Institute of Theoretical Physics and Synergetics, University of Stuttgart, April 28, 1987.

[63] Hans-Paul Schwefel. Descriptive, explicative, and normative systems analysis. First Micro-Symposium on Systems Analysis and Decision Support, University of Dortmund, Department of Computer Science, July 16, 1987.

[64] Hans-Paul Schwefel. Evolution as a paradigm of general development. IEEE Conf. Systems, Man, and Cybernetics, Alexandria VA, October 20, 1987.

[65] Hans-Paul Schwefel. Technical optimization by means of simulated evolution. 13th Int’l Conf. Mathematical Optimization – Theory and Applications, Eisenach/GDR, November 17, 1987.

[66] Hans-Paul Schwefel. Evolutionsstrategien als eine Brücke zwischen natürlicher und künstlicher Intelligenz. Karl-Marx University, Leipzig/GDR, November 20, 1987.

[67] Hans-Paul Schwefel. Evolutionsstrategien als eine Brücke zwischen natürlicher und künstlicher Intelligenz. Humboldt University at Berlin, November 22, 1987.

[68] Hans-Paul Schwefel. Technische Optimierung nach dem Vorbild der Natur – mit und ohne Computer. Hannover, VDI – Working Group on Development, Design, and Sales, Polytechnical University of Hannover, November 25, 1987.

[69] Hans-Paul Schwefel. Evolution und Optimierung: Ein Gegensatz? Second Micro-Symposium on Optimization by means of Parallel Computers, University of Dortmund, Department of Computer Science, January 12, 1988.

[70] Hans-Paul Schwefel. Künstliche Evolution als natürliche Intelligenz. German National Research Center for Information Technology (GMD), St. Augustin, January 21, 1988.

[71] Hans-Paul Schwefel. Was ist Systemanalyse, was könnte sie werden? University of Hildesheim, Institute of Computer Science, January 28, 1988.

[72] Hans-Paul Schwefel. Evolution strategies – a macroscopic view onto long-term economic processes. IFIAS-MERIT Meeting, Maastricht, The Netherlands, February 23, 1988.

[73] Hans-Paul Schwefel. Über natürliche Intelligenz und künstliche Evolution. IBM Scientific Center Heidelberg, April 27, 1988.

[74] Hans-Paul Schwefel. Problems of parameter estimation in urban system dynamics. PROCOPE Workshop on the Evolution of Urban Systems, La Petite Pierre, France, May 5, 1988.

[75] Hans-Paul Schwefel. Systemanalyse – Was ist das? University of Dortmund, Department of Spatial Planning, June 16, 1988.

[76] Hans-Paul Schwefel. Über künstliche Evolution und natürliche Intelligenz. University of Ulm, Faculty of Natural Sciences and Mathematics, June 23, 1988.

[77] Hans-Paul Schwefel. Technische versus biologische Optimierung – Kann man aus der Analogie etwas lernen? 1st Int’l Symposium of the SFB 230 Natürliche Konstruktionen – Leichtbau in Architektur und Natur, Stuttgart, September 26, 1988.

[78] Hans-Paul Schwefel. Natürliche Evolution als Vorbild für künstliche Intelligenz – ein wirtschaftlicher Weg? University of Passau, Faculty of Mathematics and Computer Science, October 21, 1988.

[79] Hans-Paul Schwefel. An evolutionary approach to multiobjective optimization. Int’l Conference Multiobjective Problems of Mathematical Programming, Yalta, USSR, October 26, 1988.

[80] Hans-Paul Schwefel. Künstliche Intelligenz und Sozioökonomie – divergente oder konvergente Entwicklungen? University of Freiburg, Faculty of Economics, Freiburg, November 8, 1988.

[81] Hans-Paul Schwefel. Optimierung nach natürlichem Vorbild. University of Kiel, Theoretical Physics Department, November 10, 1988.

[82] Hans-Paul Schwefel. Evolution strategies on multi-transputer systems. Conference Transputers in Computational Science, Center for Physicists, Bad Honnef, November 30, 1988.

[83] Hans-Paul Schwefel. Kollektives Lernen – Erfahrungen mit einem mehrgliedrigen Evolutionsmodell. Biomathematical Colloquium, University of Bonn, February 1, 1989.

[84] Hans-Paul Schwefel. Natürliche Evolution als Vorbild für künstliche Intelligenz. Research Institute for Applied Knowledge Engineering (FAW) at the University of Ulm, February 27, 1989.

[85] Hans-Paul Schwefel. Understanding evolution as a collective strategy for groping-in-the-dark. University of the Bundeswehr, Institute of Cybernetic Anthropology – Informal Workshop on Evolutionary Models and Strategies, Munich-Neubiberg, March 9, 1989.

[86] Hans-Paul Schwefel. Evolutionary development strategies in competitive environments. International Institute for Applied Systems Analysis (IIASA), Laxenburg, Austria, March 15, 1989.

[87] Hans-Paul Schwefel. Evolutionsstrategien. Academy of Sciences of the GDR, Volterra-Seminar, Berlin-Adlershof, March 28, 1989.

[88] Hans-Paul Schwefel. Über Bedingungen für kollektive Intelligenz in evolutionären Systemen. MENSA in Germany (MinD) e.V., Annual Meeting, Cologne, April 8, 1989.

[89] Hans-Paul Schwefel. Collective learning in evolutionary processes. Int’l Symposium on Evolutionary Dynamics and Nonlinear Economics, University of Texas, Austin, Ilya Prigogine Center, April 16, 1989.

[90] Hans-Paul Schwefel. Evolutionsalgorithmen. University of Dortmund, Department of Social Sciences, Philosophy, and Theology, May 2, 1989.

[91] Hans-Paul Schwefel. Evolution und Optimierung. Colloquium at the Institute of Numerical and Applied Mathematics, University of Göttingen, May 9, 1989.

[92] Hans-Paul Schwefel. Collective learning in search for extrema. Delft University of Technology, Faculty of Mathematics and Informatics, The Netherlands, June 5, 1989.

[93] Hans-Paul Schwefel. Evolution und Optimierung. Fraunhofer Institute for Production Technology and Automation, Stuttgart, August 25, 1989.

[94] Hans-Paul Schwefel. Theorie und Praxis der Evolutionsstrategien. Aachen University of Technology RWTH, Institute of Plastics Technology, September 4, 1989.

[95] Hans-Paul Schwefel. My dream of optimum-seeking. Micro-Symposium Parallel Processing and Applications, University of Siegen, November 17, 1989.

[96] Hans-Paul Schwefel. Evolutionsstrategien: Naturanaloge parallele Optimumsuche. National Research Center for Computer Sciences (GMD), University of Karlsruhe, March 7, 1990.

[97] Hans-Paul Schwefel. Evolution als Vorbild für Optimierungsstrategien. University of Freiburg, Faculty of Physics, May 4, 1990.

[98] Hans-Paul Schwefel. MCDM with evolution strategies. Int’l Workshop Adaptive Learning and Neural Networks, Schloss Reisensburg, Günzburg, June 28, 1990.

[99] Hans-Paul Schwefel. Report about the activities of the pca group at dortmund. Parallel Computing Action (PCA) Workshop, University of Southampton, July 9, 1990.

[100] Hans-Paul Schwefel. Technische Optimierung mittels Evolutionsstrategien / Evolutionsstrategien als Paradigma für stabile Entwicklungsprozesse. 25th Seminar for Technical Executives at the Aachen University of Technology RWTH, Zweifall, September 24, 1990.

[101] Hans-Paul Schwefel. Einfache Evolutionsmodelle und ihre Verwendung als Optimum-Suchalgorithmen. University of Saarbrücken, Institute for Bionics, October 8, 1990.

[102] Hans-Paul Schwefel. Conditions for synergy in collective learning (poster). Workshop MOSES – Humboldt University at Berlin, Gosen, November 5, 1990.

[103] Hans-Paul Schwefel. Technische Optimierung mittels Evolutionsstrategien. Carl-Schorlemmer University, Leuna/Merseburg, November 9, 1990.

[104] Hans-Paul Schwefel. Evolutionsstrategien: Geschichte und Perspektiven. Aachen University of Technology RWTH, Institute of Technical Thermodynamics, December 18, 1990.

[105] Hans-Paul Schwefel. Natürliche Evolution als Vorbild für den Entwurfsprozeß technischer Systeme? Design und Optimierung in Natur und Technik – Working Group Session of the Sonderforschungsbereich 230, University of Stuttgart, February 18, 1991.

[106] Hans-Paul Schwefel. Die mehrgliedrige Evolutionsstrategie als algorithmisches Modell der Höherentwicklung in natürlichen Systemen. Lotka-Volterra Seminar, University of Osnabrück, Systems Science Group, April 22, 1991.

[107] Hans-Paul Schwefel. An evolution strategy with somatic mutation for solving binary optimization problems. Workshop Combinatorial Optimization and Fitness, Interdisciplinary Research Center (IFZ), University of Bielefeld, May 6, 1991.

[108] Hans-Paul Schwefel. Evolutionäre Algorithmen für die Systemanalyse. Lecture Series Systems Engineering, University of Passau, Faculty of Mathematics and Computer Science, November 27, 1991.

[109] Hans-Paul Schwefel. Kollektive Problemlösung mittels künstlicher Evolution. Technical University of Munich, Faculty of Mathematics and Computer Science, April 10, 1992.

[110] Hans-Paul Schwefel. Über natürliche kollektive Intelligenz. Musical-Scientific Matinées of the Technical University of Ilmenau, April 26, 1992.

[111] Hans-Paul Schwefel. Evolutionäre Algorithmen in Theorie und Praxis. Siemens AG Munich, June 29, 1992.

[112] Frank Hoffmeister, Hans-Paul Schwefel, and Joachim Sprave. Parallele evolutionäre Algorithmen. Seminar of the IDDD Initiative Interdisciplinary Distributive Data processing Dortmund, University of Dortmund, July 30, 1992.

[113] Hans-Paul Schwefel. Bionik – Lohnt sich ein Neustart? Seminar Interdisciplinary Research in Architecture and Civil Engineering, University of Stuttgart, October 12, 1992.

[114] Hans-Paul Schwefel. Evolutionäre Algorithmen – naturanaloge Suche nach optimalen Lösungen. Mathematical Colloquium, Institute of Statistics and Mathematical Economics, Aachen University of Technology RWTH, January 19, 1993.

[115] Hans-Paul Schwefel. Evolutionary algorithms. Seminar Center for Economic Research, Tilburg University, The Netherlands, April 8, 1993.

[116] Hans-Paul Schwefel. Numerische Optimumsuche mittels simulierter Evolution. Max Planck Institute of Metal Research, Stuttgart, April 21, 1993.

[117] Hans-Paul Schwefel. Was können evolutionäre Algorithmen nicht? Colloquium of the Institute for Applied Computer Science, University of Wuppertal, May 11, 1993.

[118] Hans-Paul Schwefel. Experimentelle und numerische Optimierung mittels evolutionärer Algorithmen. Seminar for Technical Executives, WITCO GmbH, Bergkamen, June 16, 1993.

[119] Thomas Bäck, Frank Kursawe, and Hans-Paul Schwefel. Naturanaloge Verfahren - Grundlagen und praktischer Einsatz in der Optimierung. University of Dortmund, Tutorial of the German Academy of Computer Science, June 21, 1993.

[120] Hans-Paul Schwefel. Collective intelligence in evolving systems – some prerequisites. Symposium Artificial Life, 29th Lustrum, Delft University of Technology, The Netherlands, July 2, 1993.

[121] Hans-Paul Schwefel. Evolving evolutionary algorithms. Invited Opening Lecture, SIAM Conf. Simulation and Monte Carlo Methods, San Francisco CA, August 4, 1993.

[122] Hans-Paul Schwefel. Collective search for optimal solutions – evolutionary algorithms. Joint Seminar Series of the Applied Computing Research Institute and the Dept. of Comp. Sci. and Comp. Eng., La Trobe University, Bundoora, Victoria, Australia, September 17, 1993.

[123] Hans-Paul Schwefel. Collective search for optimal solutions – evolutionary algorithms. Seminar Series of the Machine Learning Group, Melbourne University, Carlton, Victoria, Australia, September 21, 1993.

[124] Hans-Paul Schwefel. Evolution und Optimierung – Strategien in Natur und Technik. Second FOMAAS Colloquium, Research Center for Multidisciplinary Analyses and Applied Structural Optimization, University of Siegen, December 2, 1993.

[125] Hans-Paul Schwefel. Evolutionäre Algorithmen – naturanaloge Suche nach optimalen Lösungen. Universität der Bundeswehr, Fachbereich Wirtschafts- und Organisationswissenschaften, Hamburg, February 14, 1994.

[126] Thomas Bäck, Frank Kursawe, and Hans-Paul Schwefel. Tutorial evolutionäre algorithmen. Fourth Fuzzy Days at Dortmund, June 6 – 8, 1994.

[127] Hans-Paul Schwefel. Evolutionäre Algorithmen – Wie geht es weiter? Berliner Physikalisches Kolloquium der Freien Universität, der Technischen Universität und der Humboldt Universität zu Berlin, July 14, 1994.

[128] Hans-Paul Schwefel. Was leisten evolutionäre Algorithmen? VDE/VDI-GME Congress Anwendungsfelder der Mikroelektronik, Munich, October 19, 1994.

[129] Hans-Paul Schwefel. Wann, warum und wie kann eine Heuristik mit Zufallskomponenten beim Suchen nach einer guten Entscheidung helfen? Workshop Evolutionäre Algorithmen in Management-Anwendungen, Georg-August-University, Göttingen, February 23, 1995.

[130] Hans-Paul Schwefel. Evolutionsstrategien – Künstliche Evolution als Modell für natürliche Intelligenz. Ringvorlesung Bionik an der Fachhochschule Hamburg, Fachbereich Maschinenbau und Chemieingenieurwesen, July 6, 1995.

[131] Hans-Paul Schwefel and Thomas Bäck. Tutorial on evolution strategies. Sixth Int’l Conf. Genetic Algorithms (ICGA’95), Pittsburgh PA, July 15, 1995.

[132] Hans-Paul Schwefel. Evolutionäre Algorithmen – Geschichte und Perspektiven. Schwerpunktdiskussion: Evolutionäre Algorithmen in der Automatisierungstechnik, VDI/VDE–GMA Gesellschaft für Mess- und Automatisierungstechnik, Fachausschuss 4.5 Neuronale Netze, Aachen, October 17, 1995.

[133] Hans-Paul Schwefel. What is evolutionary computation good for ? Seminar at the University of Nijmegen, Faculty of Mathematics and Computing Science, October 19, 1995.

[134] Hans-Paul Schwefel. Theoretical aspects of evolution strategies. European Short Course on Genetic Algorithms and Evolution Strategies in Computational Science and Engineering (EUROGEN’95), University of Las Palmas de Gran Canaria, Spain, December 4–8, 1995.

[135] Hans-Paul Schwefel. Introduction to evolution strategies. Rijksuniversiteit te Leiden, The Netherlands, March 8, 1996.

[136] Hans-Paul Schwefel. Missing features in current evolutionary algorithms. Dagstuhl-Seminar Evolutionary Algorithms and their Application, March 23–29, 1996.

[137] Hans-Paul Schwefel. Evolutionary computation – history, status, and perspectives. Int’l Conf. Artificial Neural Networks (ICANN’96), Ruhr-University Bochum, July 17, 1996.

[138] Hans-Paul Schwefel. Evolutionary algorithms from an evolution strategy perspective. IMA-Workshop on Evolutionary Algorithms, University of Minnesota, Minneapolis MI, October 22, 1996.

[139] Hans-Paul Schwefel. Darwin auf dem Rechner - Wozu? Swiss Federal Institute of Technology Zurich, November 1, 1996.

[140] Hans-Paul Schwefel. Computational Intelligence durch künstliche Evolution? University of Mannheim, November 18, 1996.

[141] Hans-Paul Schwefel. Vorstellung und Bericht des SFB Computational Intelligence. Fuzzy–Neuro– Initiative NRW, Duisburg, November 20, 1996.

[142] Hans-Paul Schwefel. Künstliche Evolution versus natürliche Intelligenz. 2. Dortmunder Forschungstage – Die Universität in der City, Dortmund, January 30, 1997.

[143] Hans-Paul Schwefel. Evolutionäres Rechnen in der Computational Intelligence. Fuzzy-Neuro-Systems ’97 – Computational Intelligence – Fourth Int’l Workshop, University of Paderborn, Soest, March 12, 1997.

[144] Hans-Paul Schwefel. Informatik – woher und wohin? Campusfest at the University of Dortmund, June 14, 1997.

[145] Thomas Bäck, Boris Naujoks, and Hans-Paul Schwefel. Presentation at Forschungsforum ’97 , Federal Ministry for Education, Science, Research, and Technology (BMBF), Leipzig, September 16–20 1997.

[146] Hans-Paul Schwefel. Evolutionäre Algorithmen. Ringvorlesung Innovation durch Informatik, University of Dortmund, November 24, 1997.

[147] Hans-Paul Schwefel. Evolution strategies: Basic concepts and some theory. European Short Course on Genetic Algorithms and Evolution Strategies in Computational Science and Engineering (EU-ROGEN’97), University of Trieste, Italy, November 28–December 4, 1997.

[148] Hans-Paul Schwefel. Evolutionäre Algorithmen. Technical University of Braunschweig, The Institutes of Computer Science, Colloquium, February 23, 1998.

[149] Hans-Paul Schwefel. Evolutionary algorithms – on the state of the art (keynote lecture). Third Evolutionary / Adaptive Computing in Design and Manufacture Conf. (ACDM’98), Dartington, UK, April 21, 1998.

[150] Hans-Paul Schwefel. Panel discussion. BMBF-Fachtagung Von der Natur lernen – Information in Natur und Technik, Göttingen, May 14, 1998.

[151] Hans-Paul Schwefel. Evolutionary computation – yesterday, today, and tomorrow. Leiden Center for Natural Computing, Leiden, The Netherlands, June 10, 1998.

[152] Hans-Paul Schwefel. Evolution strategies – origin, contemporary incarnations, and applications. EvoNet Summer School on Evolutionary Computation (ESSEC’98), Tours, France, July 14, 1998.

[153] Hans-Paul Schwefel. Panel discussion: The role of soft computing in the conception, design, and utilization of information/intelligent systems. EUFIT’98, Aachen, September 8, 1998.

[154] Hans-Paul Schwefel. SFB 531 design and management of complex technical processes and systems by means of computational intelligence methods. COST P4 – Int’l Workshop Non-linear dynamics in mechanical processing, Dortmund, November 6, 1998.

[155] Hans-Paul Schwefel. Evolutionary optimisation (Tutorial) / Natural computation (Keynote lecture). Second Asia-Pacific Conf. Simulated Evolution and Learning (SEAL’98), Canberra, Australia, November 24–27, 1998.

[156] Hans-Paul Schwefel. Evolutionary algorithms research at Dortmund (Seminar). Key Centre of Design Computing, University of Sydney, Australia, November 30, 1998.

[157] Hans-Paul Schwefel. Evolutionary optimisation (Tutorial). Complex Systems Summer School, Charles Sturt University, Bathurst, Australia, December 3, 1998.

[158] Hans-Paul Schwefel. Evolutionäre Algorithmen, dynamische Zielfunktionen und die Optimierung von Lichtsignalsteuerungen: Ein Überblick. HEUREKA’99 – Optimierung im Verkehr und Transport, Karlsruhe, March 4, 1999.

[159] Hans-Paul Schwefel. Evolutionsstrategien (Keynote lecture). Buchenbach-Workshop IV Evolutorische Ökonomik, May 15, 1999.

[160] Hans-Paul Schwefel. Der SFB 531 – Ziele und Stand der Forschung. Colloquium of the Three Collaborative Research Centres 376, 527, and 531, Dortmund, May 27, 1999.

[161] Hans-Paul Schwefel. Introduction to evolution strategies (tutorial). Genetic and Evolutionary Computation Conf. (GECCO 1999), Orlando FL, July 14, 1999.

[162] Hans-Paul Schwefel. Optimierung von logistischen Netzwerken (SFB 559). 17. Dortmunder Gespräche, Fraunhofer Institut für Materialfluss und Logistik, Dortmund, September 30, 1999.

[163] Hans-Paul Schwefel. Current research in evolutionary computation and computational intelligence at the university of dortmund. Evolutionary Computation Seminar, ATR Advanced Telecommunications Research Institute, Seika Science City, Japan, October 8, 1999.

[164] Hans-Paul Schwefel. Still missing features in current evolutionary algorithms. Dagstuhl-Seminar Theory of Evolutionary Algorithms, February 14, 2000.

[165] Hans-Paul Schwefel. Biologically inspired computational optimization strategies. Int’l Onsager-Workshop Optimization in Industrial and Living Systems, Oud Poelgeest, The Netherlands, March 12, 2000.

[166] Hans-Paul Schwefel. Evolutionary computation – good for what? The Hong Kong Polytechnic University, Faculty of Engineering, Dept. of Computing, Kowloon, March 27, 2000.

[167] Hans-Paul Schwefel. Classical evolutionary algorithms and their theory. Int’l Workshop on Evolutionary Computation, Wuhan University, China, April 3, 2000.

[168] Hans-Paul Schwefel. Some recent extensions of evolutionary algorithms with applications. Int’l Workshop on Evolutionary Computation, Wuhan University, China, April 5, 2000.

[169] Hans-Paul Schwefel. Evolutionary computation I: Basic concepts and some applications. University of Patras, Greece, September 26, 2000.

[170] Hans-Paul Schwefel. Evolutionary computation II: Some theory and recent advances. University of Patras, Greece, September 28, 2000.

[171] Hans-Paul Schwefel. Computación evolutiva: Panorama actual (lecture) / Mesa redonda: Computación evolutiva: una o muchas? Cursos de Verano de El Escorial: Encuentro Sobre Computación Natural, Universidad Complutense de Madrid, August 6, 2001.

[172] Hans-Paul Schwefel. Neuere Erkenntnisse über evolutionäre Algorithmen. Workshop on Biologically Inspired Methods on Modelling and Design of Circuits and Systems, Technical University Ilmenau, October 5, 2001.

[173] Hans-Paul Schwefel. Zukunft der evolutionären Algorithmen. VDI/VDE-GMA Fachausschuss 5.21 Neuronale Netze und Evolutionäre Algorithmen, Dortmund, October 19, 2001.

[174] Hans-Paul Schwefel. An introduction to evolutionary computation (tutorial). 2002 IEEE World Congress on Computational Intelligence (WCCI’02), Honolulu HI, May 12, 2002.

[175] Hans-Paul Schwefel. Was sind evolutionäre Algorithmen? Interdisciplinary workshop Computergestützte Analyse evolutionärer Optimierungsprozesse in komplexen Systemen der Berlin-Brandenburgischen Akademie der Wissenschaften, Blankensee, May 25, 2002.

[176] Hans-Paul Schwefel. Bio-inspired problem solving. Computational Intelligence in Robotics and Automation for the New Millennium (IEEE CIRA 2003), Kobe, Japan, July 17, 2003.

[177] Hans-Paul Schwefel. Genesis and future of evolutionary computing. 12th Conf. Intelligent Systems Applications in Power Systems (ISAP 2003), Lemnos, Greece, September 1, 2003.

[178] Hans-Paul Schwefel. Genesis and future of evolutionary computing. 6th Int’l Conf. Artificial Evolution (EA’03), Marseilles, France, October 28, 2003.

[179] Hans-Paul Schwefel. Evolutionäre Struktur- und Parameteroptimierung verfahrenstechnischer Prozesse nach mehreren Kriterien. 2nd Workshop of the BBAW Study Group Structure Evolution and Innovation, Berlin-Brandenburgische Akademie der Wissenschaften (BBAW), Berlin, November 21, 2003.

[180] Hans-Paul Schwefel. A private view onto evolutionary computation: Its past, present, and future. School of Computer Science, The University of Birmingham, UK, September 17, 2004.

[181] Hans-Paul Schwefel. Natural computing – why and how? Indian Institute of Technology, Kanpur, India, October 1, 2004.

[182] Hans-Paul Schwefel. Does evolutionary computation have a future? The Hong Kong Polytechnic University, Faculty of Engineering, Dept. of Computing, Kowloon, March 11, 2005.

[183] Hans-Paul Schwefel. Evolutionary optimization: What I have learned since 1964 – Part I and II. ISICA 2005, China University of Geosciences, Wuhan, April 4/5, 2005.

[184] Peter Schuster, Klaus Lucas, Peter Roosen, and Hans-Paul Schwefel. Evolutionär ablaufende Strukturbildung. Ratspräsentation, Berlin-Brandenburgische Akademie der Wissenschaften (BBAW), Berlin, May 26, 2005.

[185] Hans-Paul Schwefel. Closing words: Life is a game of dice – La vie est un jeu de dé. Colloquium on the Occasion of the Advent of his Retirement from the Dept. of Computer Science, University of Dortmund, February 11, 2006.

[186] Günter Rudolph on behalf of Hans–Paul Schwefel. Acknowledgment for the elevation to IEEE Fellow. Awards Banquet, 2006 IEEE World Congress on Computational Intelligence (WCCI’06), Vancouver, Canada, July 19, 2006.

[187] Hans-Paul Schwefel. Acknowledgment for the admission to the degree of doctor of science honoris causa. Degree Congregation, The University of Birmingham, Birmingham, UK, December 13, 2007.

[188] Hans-Paul Schwefel. A pioneer’s view onto evolutionary optimisation (invited opening talk). EvoStar 2008 Conferences, Napoli, March 26, 2008.

[189] Hans-Paul Schwefel. Simulated evolution under multiple criteria conditions revisited (invited plenary talk). 2008 IEEE Congress on Evolutionary Computation (CEC) as part of the 2008 IEEE World Congress on Computational Intelligence (WCCI’08), Hong Kong, June 3, 2008.

[190] Hans-Paul Schwefel. Panel discussion: Challenges and opportunities of computational intelligence research – Pioneers speaking. 2008 IEEE Congress on Evolutionary Computation (CEC) as part of the 2008 IEEE World Congress on Computational Intelligence (WCCI’08), Hong Kong, June 4, 2008.

[191] Hans-Paul Schwefel. Mutation, recombination, and selection revisited (pioneer lecture 1). Workshop and Summer School on Evolutionary Computing and Lecture Series by Pioneers, University of Ulster Magee Campus, Intelligent Systems Research Centre and School of Computing and Intelligent Systems, Derry/Londonderry, Northern Ireland, UK, August 18, 2008.

[192] Hans-Paul Schwefel. Beyond mutation, recombination, and selection (workshop lecture 1). Workshop and Summer School on Evolutionary Computing and Lecture Series by Pioneers, University of Ulster Magee Campus, Intelligent Systems Research Centre and School of Computing and Intelligent Systems, Derry/Londonderry, Northern Ireland, UK, August 18, 2008.

[193] Hans-Paul Schwefel. Panel discussion. Workshop and Summer School on Evolutionary Computing and Lecture Series by Pioneers, University of Ulster Magee Campus, Intelligent Systems Research Centre and School of Computing and Intelligent Systems, Derry/Londonderry, Northern Ireland, UK, August 20, 2008.

[194] Hans-Paul Schwefel. Jubilee talk (opening address). 10th International Conference on Parallel Problem Solving from Nature (PPSN X), Dortmund, September 15, 2008.

[195] Hans-Paul Schwefel. Nicht-Standard Evolutionäre Algorithmen. University of Rostock, Faculty of Computer Science and Electrical Engineering, October 16, 2008.

[196] Hans-Paul Schwefel. Opening address. WSC 2008 Online World Conference on Soft Computing in Industrial Applications, November 10-28, 2008.

[197] Hans-Paul Schwefel. Simulierte Evolution - gestern und morgen. 18th Workshop Computational Intelligence of the GMA-FA 5.14 Computational Intelligence and the GI-FG on Fuzzy Systems and Soft Computing, Universitätskolleg Bommerholz of Dortmund University of Technology at Witten-Bommerholz, December 5, 2008.

[198] Hans-Paul Schwefel. Was sind und was leisten Evolutionsstrategien? Colloquium Biology and Society, Interdisciplinary series of events of the section Biology, Dortmund University of Technology, January 12, 2009.

[199] Hans-Paul Schwefel. Failures as stepping stones to success or per aspera ad astra (invited talk). Workshop Learning from Failures in Evolutionary Computation, 2009 Genetic and Evolutionary Computation Conference (GECCO’09), Montréal, Canada, July 8, 2009.

[200] Hans-Paul Schwefel. eVolution seit dem Beginn der Informatik (invited talk). 39th Annual Conference of the Gesellschaft für Informatik (GI), Workshop Mit Leben rechnen – Zur Geschichte des Wissenstransfers zwischen Computer- und Biowissenschaften of the Special Interest Group on Informatics and Computing History, Lubeck, September 29, 2009.

[201] Hans-Paul Schwefel. Zurück in die Zukunft – Ein Zeitzeuge berichtet (invited testimonial lecture). Jubiläums-Veranstaltung 10 Jahre IuK-Ausschuss, Chamber of Industry and Commerce at Dortmund, November 10, 2009.

[202] Hans-Paul Schwefel. Darwin auf dem Computer – Innovation durch simulierte Evolution. FORUM Der strategische Mensch: Zur Evolution von Strukturen im Rahmen des Jahresthemas 2009/2010 Evolution in Natur, Technik und Kultur of the Berlin Brandenburg Academy of Sciences and Humanities, Berlin, February 2, 2010.

[203] Hans-Paul Schwefel. Chancen und Risiken des Zufalls in evolutionären Algorithmen. MENSA in Deutschland (MinD) e.V., Annual Meeting, Dortmund, Symposium Zufall, Risiko und Chancen, April 16, 2010.

[204] Hans-Paul Schwefel. Interview on the occasion to celebrate Konrad Zuse’s 100th birthday. Radio Dortmund 91,2 MHz, June 22, 2010.

[205] Hans-Paul Schwefel. On forgotten island treasures and future jungle paths (invited opening lecture). 11th International Conference on Parallel Problem Solving from Nature (PPSN XI), Cracow, Poland, September 13, 2010.

[206] Hans-Paul Schwefel. 30 Jahre Informatik Lehrstuhl XI (internal presentation). Chair of Algorithm Engineering, Dept. of Computer Science, University of Dortmund, December 15, 2015.

[207] Hans-Paul Schwefel. 30 years ppsn ... from so simple a beginning (invited opening lecture). 16th International Conference on Parallel Problem Solving from Nature (PPSN XVI), Leiden, The Netherlands, September 5, 2020.